The modified Phillips curve is a relation between the change in the inflation rate and the level of the unemployment rate. Explain the logic of this curve and how it differs from the original Phillips curve. Using this Phillips curve, is the unemployment rate zero when the rate of inflation is neither rising nor falling? Why do Central Banks focus so much on whether inflation expectations are well anchored? Explain.        



1
Expert's answer
2021-02-15T08:59:20-0500

Philips curve shows the trade-off between the inflation rate and the unemployment and depicts the negative relationship between both.

Original Philips Curve Equation:

"\pi t = \bar{\pi}+(m+z) -au_t"

Modified Philips curve Equation:

"\pi_t - \pi_{t-1} = (m+z) -au_t"

The original Philips curve was based on the assumption that expected inflation does not change over time and remains fixed. So, "\bar{\pi} = \pi _{t}^{e}" . The equation implies that policymakers can tolerate higher inflation by maintaining a lower unemployment rate forever, so there exists no natural rate of unemployment and the trade-off between inflation and unemployment could exist only if wage setters systematically underpredict inflation and make some mistakes forever and if the government attempts to sustain lower unemployment by accepting higher unemployment, the trade-off would ultimately disappear.

But in the 1970s (oil crisis), this trade-off is broken down. The Philips curve had to be modified. So, the modification was done in the following way:

"\pi _{t}^{e} = (1-\theta )\bar{\pi } + \theta \pi _{t-1}"

Modified Philips curve becomes: "\pi t = \theta \pi _{t-1}+(m+z)-\alpha u_{t}"

if "\theta = 0" , then the above equation becomes the original Philips curve

if "\theta" is positive, then actual inflation is dependent upon last year's inflation

The above equation is also known as "Expectations Augmented Philips Curve"

Change in inflation depends upon the difference between the actual and natural unemployment rate. The natural rate of unemployment is that rate that is required to keep inflation constant. When the inflation rate is neither falling nor increasing, the natural rate of unemployment is not zero.

Central Banks put much focus on anchoring the expected inflation because expected inflation paves the way for how people perceive how the economy is going to be in the future. Anchoring of expected inflation implies a steady inflation rate in the future which also creates the credibility of the Central Bank Policies among the general public. It gives the indication that Central Bank is determined to control the inflation rate through its monetary policy tools.